OpenStudy (anonymous):

Determine the number of possible solutions if A=62degrees, a=4, and b=6

OpenStudy (dumbcow):

use law of sines to find angle B

OpenStudy (anonymous):

would you be able to please walk me through that?

OpenStudy (anonymous):

wait hold on

OpenStudy (dumbcow):

\[\frac{sinB}{6} = \frac{\sin 62}{4}\]

OpenStudy (anonymous):

ok and then 1.613/25 for C and c

OpenStudy (dumbcow):

you will get 2 possible answers for B

OpenStudy (anonymous):

but how does that determine the # of possible solutions?

OpenStudy (dumbcow):

since you can have 2 answers for B, that leads to 2 answers for c and C as well hence 2 possible triangles

OpenStudy (dumbcow):

whenever you take the inverse sin to solve you get 2 solutions this comes from unit circle